OpenStudy (anonymous):

derive:dR/dT = d/dT ((R1*R2)/(R2+R1))

OpenStudy (anonymous):

\[\frac{ dR }{ dT } = \frac{ d }{ dT } \left( \frac{ R_1.R_2 }{ R_2 + R_1 } \right)\]

OpenStudy (anonymous):

do i derive the num and denom separately?

OpenStudy (anonymous):

You don't "Derive"; You "Differentiate". Now moving on...

OpenStudy (anonymous):

@Rowa So what are you asked to do in this problem?

OpenStudy (anonymous):

it's alot to explain, but it comes down to the equation i gave

OpenStudy (anonymous):

I know but what are we supposed to do with the equation? @Rowa

OpenStudy (anonymous):

just says, calculate dR/dT

OpenStudy (anonymous):

and R=R_1*R_2/R_2+R_1

OpenStudy (anonymous):

Oh ok. So Are \(\bf R_1 \) and \(\bf R_2\) functions or constants? @rowa

OpenStudy (anonymous):

functions

OpenStudy (anonymous):

What are they functions of? Like are they both functions of 'x' or what? @Rowa

OpenStudy (anonymous):

of T

OpenStudy (anonymous):

OOOOOOOOOOOH. So basically, this is what it is:\[\bf R(T)=\frac{ R_1(T) \frac{}{} R_2(T) }{ R_1(T)+R_2(T) }\] @Rowa Correct?

OpenStudy (anonymous):

guess so

OpenStudy (anonymous):

Ok so then we basically use the Division rule to differentiate:\[\bf \frac{ dR }{ dT }=\frac{ (R_1 \frac{}{} R_2)'(R_1(T)+R_2(T))-(R_1(T)+R_2(T))'(R_1 \frac{}{} R_2)}{ (R_1(T)+R_2(T))^2 }\]

OpenStudy (anonymous):

done, i was doing that while waiting

OpenStudy (anonymous):

and why is it differentiating?

OpenStudy (anonymous):

ya it takes a while to type out lol

OpenStudy (anonymous):

@Rowa When we take the "Derivative of something like a function", we are differentiating the function not deriving it. This is a common misconception.

OpenStudy (anonymous):

alright, it's just a way of saying it

OpenStudy (anonymous):

@Rowa Have you simplified the top and bottom yet?

OpenStudy (anonymous):

Leave the denominator for last, finish up with the top first.

OpenStudy (anonymous):

i applied the product rule and addition rule

OpenStudy (anonymous):

Ya exactly. The product rule and the addition rule and then you're done. You can simplify further after that but I believe there won't be much simplification in a situation like this. So after applying that, you should be fine, assuming you applied it correctly. @Rowa

OpenStudy (anonymous):

maybe i should type the complete question, because i'm not understanding some steps written by my professor

OpenStudy (anonymous):

somehow, R_1 dR_2/dT is zero..

OpenStudy (anonymous):

I'm pretty sure that this is not good for me because I haven't seen the whole question so I can't really say anything. The best thing to do is to take a picture of the question and then attach it and post.

OpenStudy (anonymous):

problem is, it's in a different language

OpenStudy (anonymous):

i'll have to translate it

OpenStudy (anonymous):

Where are you from lol

OpenStudy (anonymous):

europe

OpenStudy (anonymous):

.......Europe isn't the greatest answer..........

OpenStudy (anonymous):

western europe?

OpenStudy (anonymous):

isn't there a country you live in?

OpenStudy (anonymous):

well yea, but is that important?

OpenStudy (anonymous):

shall i translate?

OpenStudy (anonymous):

ok just post the whole question translated. that would be best. and post the original question alongside as well.

OpenStudy (anonymous):

2 resistors R_1 and R_2 are parallel to each other. The total R is 1/R = 1/R_1 + 1/R_2 R_1 is heated slowly, this causes an increase of the resistor given by : R_1(T) = R_0 (1+T-T_0/300) T is temperature, R_0 and T_0 are constants that givethe resistance and temperature of R_1 at time 0. The question: Give the instantaneous increase in R (in function of T), so : dR/dT

OpenStudy (agent0smith):

R2 is a constant then, not a function. Makes this problem simpler.

OpenStudy (anonymous):

how is R_2 a constant

OpenStudy (agent0smith):

Only R1 is heated. R2's resistance is constant.

OpenStudy (anonymous):

@Rowa Aren't you given what R2 is? Or is R2 defined as a function of T as well?

OpenStudy (anonymous):

don't know what R2 is, i've typed everything it says here

OpenStudy (anonymous):

@agent0smith R2 isn't the same. \[\bf \frac{ 1 }{ R }=\frac{ 1 }{ R_1 }+\frac{ 1 }{ R_2 }\]Since R1 is a function then 1/R2 varies as R1 changes hence R2 is not constant.

OpenStudy (anonymous):

sure? in the steps, it says the derivative of R2 is zero

OpenStudy (anonymous):

Then R2 is a constant!

OpenStudy (agent0smith):

R2 is a constant... A resistors resistance can't change all on it's own.

OpenStudy (anonymous):

hmm, i see

OpenStudy (anonymous):

@Rowa You gotta tell the whole story or we won't know otherwise lol.

OpenStudy (agent0smith):

@genius12 1/R varies as 1/R1 varies. 1/R2 is just a constant.

OpenStudy (anonymous):

@agent0smith ya ya I realise lol.

OpenStudy (agent0smith):

Total resistance R is NOT constant.

OpenStudy (anonymous):

alright, but the steps i took to differentiate are correct?

OpenStudy (anonymous):

comparing them to the steps my prof. took, they look correct

OpenStudy (anonymous):

Ok, now going back to what you said: \[\bf R=\frac{ R_1 \frac{}{} R_2 }{ R_1 + R_2 }\]Then the derivative is evaluated the same way as we did with the division rule except now wherever you see R1, replace it with the function R1 given in full form and consider R2 to be a constant and simplify from there on. @Rowa

OpenStudy (anonymous):

\[\frac{ R_2+R_1(R_1+R_2\frac{ dR_1 }{ dT }) -(R_1R_2) + \frac{ dR_1 }{ dT } }{ (R_2+R_1)^2 }\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@Rowa Scroll up to the division rule derivative that I gave you. Just replace R1 with the full form of the function given and consider R2, R0, T0 as constants. The only variable will be T. And simplify from there.

OpenStudy (anonymous):

is what i wrote incorrect?

OpenStudy (anonymous):

@Rowa Yes.

OpenStudy (anonymous):

forgot to remove the R_1 there

OpenStudy (anonymous):

\[\frac{ R_2+R_1(R_2\frac{ dR_1 }{ dT }) -(R_1R_2) + \frac{ dR_1 }{ dT } }{ (R_2+R_1)^2 }\]

OpenStudy (anonymous):

@Rowa Now that we know that R2 is a constant, it should be:\[\bf \frac{ dR }{ dt }=\frac{ \left( R_2\frac{ dR_1 }{ dT }\right)(R_1(T)+R_2)-\left( \frac{ dR_1 }{ dT } \right)(R_1(T) \frac{}{} R_2) }{ (R_1(T)+R_2)^2 }\]

OpenStudy (anonymous):

put a plus there by accident, but other than that..my equation is good..?

OpenStudy (anonymous):

\[\frac{ R_2+R_1(R_2\frac{ dR_1 }{ dT }) -(R_1R_2) \frac{ dR_1 }{ dT } }{ (R_2+R_1)^2 }\]

OpenStudy (anonymous):

Yup that's good.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

now to replace R1 with R_0 (1+T-T_0/3000)

OpenStudy (anonymous):

Now, is R1 this:\[\bf R_1(T)=R_0\left( 1+T-\frac{ T_0 }{ 300 } \right)\]? @Rowa

OpenStudy (anonymous):

\[\bf R_1(T)=R_0\left( 1+\frac{ T-T_0 }{ 3000 } \right)\]

OpenStudy (anonymous):

Then:\[\bf \frac{ dR_1 }{ dT }=\frac{ R_0 }{ 3000 }\]Since \(\bf R_0, T_0\) are constants. @Rowa

OpenStudy (anonymous):

So now you can replace dR/dt with R0/3000 and replace R1 with the full form. If simplification is possible, simplify and you're done.

OpenStudy (anonymous):

i see, so: \[\frac{ dR_0 }{ dT } +\frac{ d }{ dT }(\frac{ R_0T-T_0 }{ 3000})\]

OpenStudy (anonymous):

@Rowa I just realised the R2 + R1 should be in brackets in the numerator of your derivative.

OpenStudy (anonymous):

dR_0 is zero, T_0 is zero, leaves me with R_0T/3000 = R_0/3000

OpenStudy (anonymous):

oh yes, forgot

OpenStudy (anonymous):

So dR_1/dT = R_0/3000

OpenStudy (anonymous):

just writing the steps i take, to see if its good

OpenStudy (anonymous):

Can you carry on from here or do you still need me?

OpenStudy (anonymous):

i think i can, thanks for helping!

OpenStudy (anonymous):

the answer should be \[\frac{ R_2^2R_0 }{ 3000(R_1+R_2)^2 }\]

OpenStudy (anonymous):

but i get \[\frac{ R_2^2\frac{ R_0 }{ 3000 } }{ (R_1+R_2)^2 }\]

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@genius12 could you verify?

OpenStudy (anonymous):

@rowa Both answers are the same thing. You see this when you write your answer like this:\[\bf \frac{ R_2^2 \frac{ R_0 }{ 3000 } }{ (R_1+R_2)^2 }=\frac{ 1 }{ (R_!+R_2)^2 } \frac{}{} R_2^2\frac{ R_0 }{ 3000 }=\frac{ R_2^2R_0 }{ 3000(R_1+R_2)^2 }\]So you see, they're basically the same answer written in differently. Hope that helps.